OpenStudy (anonymous):

I need to write cos^2(x)+2cos(x)+1 in factored form including only one function. The answer is (cos(x)+1)^2 but how would you get to that answer?

OpenStudy (anonymous):

\[ \begin{split} &\cos^2(x)+2\cos(x)+1\\ =&\cos^2(x)+\cos(x)+\cos(x)+1\\ =&\cos(x)(\cos(x)+1)+1(\cos(x)+1)\\ =&(\cos(x)+1)^2 \end{split} \]

OpenStudy (anonymous):

So one way, is splitting the middle.

OpenStudy (anonymous):

Let's think of cos in terms of x for a second. Instead of: \[\cos ^{2}x+2cosx+1\] We could write a similar equation: \[x ^{2}+2x+1\] When we factor this we get: (x+ )(x+ ) a=1 times c=1 gives us 1. Factors of one that equal two are one and one (woah). so, (x+1)(x+1) is the factored form of: \[x ^{2}+2x+1\] Now, for \[\cos ^{2}x+2cosx+1\] we follow the same principles. (cosx+ )(cosx+ ) (cosx+1)(cosx+1) or (cosx+1)^2

OpenStudy (anonymous):

Is this making sense now?

OpenStudy (anonymous):

Yes thank you so much. I tried to get help from my parents, but I guess I've surpassed them in their mathematical knowledge.

OpenStudy (anonymous):

One question of you're still there. Does this strategy work once other functions are thrown into the mix, or would I have to do something else?

OpenStudy (anonymous):

It works best if you have all the same function.

OpenStudy (anonymous):

Okay thanks.

OpenStudy (anonymous):

Usually you can use trigonometric identities to perform substitutions. If that's the case, go for it. You might be able to cancel out a function and put it in terms of a single function. Just be suspicious of 1-2sinx, 1, cos^2x + sin^2x, 1-sin^2x, 1-cos^2x. The same goes for the other pythagorean identities.

OpenStudy (anonymous):

You know, tan^2x+1=sec^2x and 1+cot^2x=csc^2x

OpenStudy (anonymous):

Yeah:) Thank you so much, this was a life saver.

OpenStudy (anonymous):

Glad to hear that :D If you have any more trig questions then feel free to ask!